# Finite Index and Do Carmo’s Question for Constant Mean Curvature Hypersurfaces

### v1

@@ Abstract
We prove that any finite $\delta$-index hypersurface $M$ in ${\mathbb{R}}^{n+1}$ with constant mean curvature must be minimal, provided
@@
  • the Ricci curvature of $M$ satisfies $\operatorname{Ric}_{M}\geq-\frac{4(1-\delta)}{n-1}|A|^{2}g,$ where $A$ is the second fundamental form and $g$ is the metric on $M.$
@@
Notice that, we do not have any restriction on the dimension and that the second result is new also in the case of finite index hypersurfaces ($\delta=0$).
@@ Introduction
Is a noncompact, complete, stable, constant mean curvature hypersurface of ${\mathbb{R}}^{n+1}$ necessarily minimal ($H=0$)?
@@
Do Carmo’s question has been investigated in ${\mathbb{R}}^{n+1}$ and in more general ambient manifolds and in the broader class of finite index constant mean curvature hyperurfaces. A positive answer for $n=3,4$ was obtained independently by [ENR] and Cheng [Che], using the Bonnet-Myers method nowadays also known as the conformal method. More precisely, there is no stable constant mean curvature hypersurface in a manifold $N$ of dimension $n+1,$ $n=3,4$ provided $H$ is large enough with respect to the sectional curvature of the ambient manifold $N$. In particular, in the Euclidean space this implies nonexistence for any $H\not=0$. Moreover, partial results for a hypersurface $M$ of arbitrary dimension have been obtained by several authors (see for example [AdC] for $M$ with polynomial volume growth , [INS2] for $M$ with subexponential volume growth and [AlCa, DCZh] for $M$ satisfying assumptions on the total curvature, and $n\leq 5$).
@@
A complete, non-compact, finite $\delta$-index, $H$-hypersurface $M$ with $\operatorname{Ric}_{M}\geq-\frac{4(1-\delta)}{n-1}|A|^{2}g,$ immersed ${\mathbb{R}}^{n+1}$ is a hyperplane.
@@ 1 Definitions and Basic Properties
Throughout this article $M$ will be a complete, noncompact $n$-dimensional hypersurface with constant mean curvature $H$ ($H$-hypersurface) of a Riemannian manifold $N$.
@@ 3 Ricci curvature bounded below
$\mathrm{Ric}_{M}>-\frac{4(1-\delta)}{n-1}|A|^{2}g.$
@@
$\operatorname{Ric}_{M}\geq-\frac{4(1-\delta)}{n-1}|A|^{2}g.$
@@
In the following subsection, we recall some useful definitions and results from criticality theory, referring the reader to [CMMR] and [PRS] for a comprehensive exposition.
@@ 3.2 The proof of the theorem
$\begin{split}-k\int_{0}^{+\infty}\varphi^{2}\frac{\hat{L}_{V}u}{u}\,ds&+\int_{0}^{+\infty}\varphi^{2}\left(R_{11}+k\hat{V}\right)\,ds+k\int_{0}^{+\infty}\frac{|\nabla u|^{2}}{u^{2}}\leq\\ &\leq(n-1)\int_{0}^{+\infty}(\varphi_{s})^{2}\,ds+k(n-1)\int_{0}^{+\infty}\varphi\varphi_{s}u_{s}u^{-1}\,ds+\frac{k^{2}(n-1)}{4}\int_{0}^{+\infty}\varphi^{2}u_{s}^{2}u^{-2}\,ds\\ &-\int_{0}^{+\infty}\varphi^{2}k(n-2)(\log u)_{ss}\,ds.\end{split}$
@@
$k\int_{0}^{+\infty}\frac{-\hat{L}_{V}u}{u}\varphi^{2}\leq\left\{k\left[\frac{(n-1)k}{4}-1\right]+\varepsilon\right\}\int_{0}^{+\infty}\varphi^{2}(u_{s})^{2}{u^{-2}}+\left\{(n-1)+\frac{k^{2}(n-3)^{2}}{4\varepsilon}\right\}\int_{0}^{+\infty}(\varphi_{s})^{2},$
@@
Now, take $0\leq\varphi_{R}\leq 1$ such that $\varphi_{R}\equiv 1$ on $[a,R]$, for some fixed $a$, $\varphi_{R}=0$ on $[R,2R]$ and $|(\varphi_{R})_{s}|\leq c/R$, where $c$ is a constant depending on $a.$ Letting $R\to\infty$, one obtains $\hat{L}_{V}u=0$ on $\gamma$. Since the above argument only uses the fact that $u$ is a positive supersolution of $\hat{L}$, we conclude that every positive supersolution $u$ of $\hat{L}_{V}$ satisfies
@@
Assume by contradiction that $\hat{L}_{V}$ is subcritical. By Theorem 3.5 there exists a positive function $W\in C(M\setminus K)$ such that $\hat{L}_{V}+W\geq 0$. By approximating $W$ from below, we can assume $W\in C^{\infty}(M\setminus K)$. Then, there exists a positive $u\in C^{2,\alpha}_{loc}(M\setminus K)$ such that $L_{V}u+Wu\leq 0$, that is $\hat{L}_{V}u\leq-Wu<0$ on the entire $M\setminus K$ and this is impossible.
@@ Appendix A Caccioppoli’s type inequalities
The purpose of this section is to prove some Caccioppoli’s type inequalities for $\delta$-stable $H$-hypersurfaces.
Let $N$ be an orientable Riemannian manifold and let $M$ be a noncompact, orientable hypersurface immersed in $N$ with constant mean curvature.
When the mean curvature is nonzero, we orient $M$ by its mean curvature vector $\vec{H}$. Then $\vec{H}=H\nu$, with $H>0$ and $\nu$ a unit normal vector field.
When the mean curvature is zero, we choose once and for all an orientation $\nu$ on $M$. We denote by $\phi$ the traceless part of the second fundamental form, i.e. $\phi:=A-Hg$, where $g$ is the induced metric on $M$.
We denote by $\nabla$ the Levi-Civita connection on $N$, and by $R$ (respectively, $\operatorname{Ric}$) the Riemann curvature tensor (respectively, the Ricci curvature) of $N$. We denote by $\nabla$ also the induced connection on $M$ for the induced metric $g$.
@@
Furthermore, we assume that the derivative of the curvature tensor is bounded. That is, there exists a constant $K^{\prime}$ such that, for any elements $e_{i},e_{j},e_{k},e_{s},e_{t}$ of a local orthonormal frame, one has
@@
The following Simon’s type inequality for H-hypersurfaces will be useful.
Theorem A.1 (Simon’s type inequality([INS1], Theorem 3.1)).
Let $M$ be a hypersurface with constant mean curvature $H$, immersed in a Riemannian manifold $N$.
Then, there exists $\varepsilon>0$ such that the following Simons-type inequality holds:
$\begin{split}|\phi|\,\Delta|\phi|\geq&\frac{2}{n(1+\varepsilon)}|\nabla|\phi||^{2}-|\phi|^{4}-\frac{n(n-2)H}{\sqrt{n(n-1)}}|\phi|^{3}+nH^{2}|\phi|^{2}+\frac{(K_{2}-K_{1})H}{2}|\phi|^{3}\\ &+(2K_{2}-K_{1})|\phi|^{2}-2nK^{\prime}|\phi|+\left(\frac{n^{2}H(K_{2}-K_{1})}{2}-\frac{n(n-1)}{2\varepsilon}(K_{1}-K_{2})^{2}\right).\end{split}$
@@
and
$a_{1}$ $=\frac{n(n-2)}{\sqrt{n(n-1)}},$
$a_{2}$ $=\frac{n}{2}(K_{2}-K_{1})H+n(2K_{2}-K_{1}),$
@@
Proof.
Let $K\subset M$ be such that $M\setminus K$ is $\delta$- stable. Inequality (A) in this case (with $\varepsilon=0,K=a_{3}=0,a_{2}=nc,q\geq 0$) yields:
$\left(\frac{2}{n}+(2q+1)-\tilde{\varepsilon}\right)\int_{M\setminus K}|\phi|^{2q}f^{2}|\nabla|\phi||^{2}\leq\frac{1}{\tilde{\varepsilon}}\int_{M\setminus K}|\nabla f|^{2}|\phi|^{2q+2}+\int_{M\setminus K}|\phi|^{2q+4}f^{2}$
$+a_{1}H\int_{M\setminus K}|\phi|^{2q+3}f^{2}-n(H^{2}+c)\int_{M\setminus K}|\phi|^{2q+2}f^{2}.$
From Young’s inequality, for any $\eta>0$,
$|\phi|^{2q+3}f^{2}\leq\frac{\eta}{2}|\phi|^{2q+4}f^{2}+\frac{1}{2\eta}|\phi|^{2q+2}f^{2}.$
and then,
$\left(\frac{2}{n}+(2q+1)-\tilde{\varepsilon}\right)\int_{M\setminus K}|\phi|^{2q}f^{2}|\nabla|\phi||^{2}\leq\frac{1}{\tilde{\varepsilon}}\int_{M\setminus K}|\nabla f|^{2}|\phi|^{2q+2}$ (7)
$+\left(1+\frac{a_{1}H\eta}{2}\right)\int_{M\setminus K}|\phi|^{2q+4}f^{2}+\left(\frac{a_{1}H}{2\eta}-n(H^{2}+c)\right)\int_{M\setminus K}|\phi|^{2q+2}f^{2}.$
Moreover, inequality (A) in this case (with $K_{2}=c$) gives:
$-(q+1)(q+1+\tilde{\varepsilon})\int_{M\setminus K}|\phi|^{2q}f^{2}|\nabla|\phi||^{2}$ $\leq\left(1+\frac{q+1}{\tilde{\varepsilon}}\right)\int_{M\setminus K}|\nabla f|^{2}|\phi|^{2q+2}$ (8)
$-\int_{M\setminus K}f^{2}|\phi|^{2q+2}(1-\delta)(|\phi|^{2}+n(H^{2}+c)).$
[…]
After rearranging terms, we obtain:
$P_{\delta}\int_{M\setminus K}|\phi|^{2q+4}f^{2}\leq L\int_{M\setminus K}|\phi|^{2q+2}|\nabla f|^{2}+Q_{\delta}\int_{M\setminus K}|\phi|^{2q+2}f^{2},$ (9)
where:
$P_{\delta}$ $=\left(\frac{2}{n}+2q+1-\tilde{\varepsilon}\right)(1-\delta)-(q+1)(q+1+\tilde{\varepsilon})\cdot\left(1+\frac{a_{1}H\eta}{2}\right),$
[…]
$Q_{\delta}$ $=(q+1)(q+1+\tilde{\varepsilon})\cdot\left(\frac{a_{1}H}{2\eta}-n(H^{2}+c)\right)-\left(\frac{2}{n}+2q+1-\tilde{\varepsilon}\right)n(H^{2}+c)(1-\delta).$
Fix $R_{0}$ such that $K\subset B^{M}(R_{0})$, so $M\setminus B^{M}(R_{0})$ is $\delta$-stable. Define $f\in C_{0}^{\infty}(M\setminus B^{M}(R_{0}))$ to be a radial function such that
$f\equiv 0\text{ on }B^{M}(R),\quad f\equiv 1\text{ on }B^{M}(R_{0}+R+1)\setminus B^{M}(R_{0}),\quad|\nabla f|\leq C.$
Replacing this $f$ in (9) gives:
$\int_{B^{M}(R_{0}+2R+1)\setminus B^{M}(R_{0})}|\phi|^{2q+4}\leq S_{\delta}\int_{M\setminus B^{M}(R_{0})}|\phi|^{2q+2},$
where $S_{\delta}=\frac{L+C^{2}Q_{\delta}}{P_{\delta}}$. Letting $R\to\infty$ we conclude
$\int_{M\setminus B^{M}(R_{0})}|\phi|^{2q+4}\leq S_{\delta}\int_{M\setminus B^{M}(R_{0})}|\phi|^{2q+2}.$
Note that the positivity of $P_{\delta}$ is equivalent to the positivity of
$\left(\frac{2}{n}+2q+1\right)(1-\delta)-(q+1)^{2}$
and so $S_{\delta}$ is positive for $0\leq q<-\delta+\sqrt{\delta^{2}-\delta+\frac{2}{n}(1-\delta)}$.
Observe that, in order to have some $q$ satisfying this inequality
$\delta+\sqrt{\delta^{2}-\delta+\frac{2}{n}(1-\delta)}>0,$
that implies $\delta<\frac{2}{n+2}$. ∎
@@
    c=1c=-1, $\varepsilon>0$, $q\in[1,\operatorname{min}\{\alpha_{2},q_{2}\}-\varepsilon]$, and $H^{2}\geq g_{n}(q)$.
@@
    The only case with no condition on $|\phi|$ is $A_{\delta}>0$ and $\Delta_{-1}<0$. One needs $q\in(\operatorname{max}\{\alpha_{1},q_{1}\},\operatorname{min}\{\alpha_{2},q_{2}\})$ and $H^{2}>g_{n}(q)$, where $g_{n}$ is defined as
@@
Proof.
Let $K$ be the compact set in $M$ such that $M\setminus K$ is $\delta$- stable. From Theorem A.2 (rearranging terms) we have
$\int_{(M\setminus K)^{+}}A_{\delta}f^{2}|\phi|^{2q+4}\leq D\int_{(M\setminus K)^{+}}|\phi|^{2q+2}|\nabla f|^{2}-BH\int_{(M\setminus K)^{+}}f^{2}|\phi|^{2q+3}$
$-(E_{\delta}+C_{\delta}H^{2})\int_{(M\setminus K)^{+}}f^{2}|\phi|^{2q+2}+F\int_{(M\setminus K)^{+}}f^{2}|\phi|^{2q+1}+G\int_{(M\setminus K)^{+}}f^{2}|\phi|^{2q}$
Using Young’s inequality, we transform all the terms on the right hand side to obtain terms involving $f^{2}|\phi|^{2q+4}$:
$|B|H|\phi|^{2q+3}$ $\leq\varepsilon_{1}|\phi|^{2q+4}+\frac{1}{\varepsilon_{1}}(|B|H)^{2}|\phi|^{2q+2}$
$|E+CH^{2}||\phi|^{2q+2}$ $\leq\varepsilon_{2}|\phi|^{2q+4}+\frac{1}{\varepsilon_{2}}|E+CH^{2}||\phi|^{2q}$
$F|\phi|^{2q+1}$ $\leq\varepsilon_{3}|\phi|^{2q+4}+\frac{1}{\varepsilon_{3}}F^{2q+4/3}$
$G|\phi|^{2q}$ $\leq\varepsilon_{4}|\phi|^{2q+4}+\frac{1}{\varepsilon_{4}}|G|^{q+2}$
for any $\varepsilon_{i}$, $i=1,2,3,4$.
Define $M^{\prime}=\{p\in M\mid f(p)\neq 0\}$. Then, on $M^{\prime}$,
$|\phi|^{2q+2}|\nabla f|^{2}=f^{2}\cdot\frac{|\phi|^{2q+2}|\nabla f|^{2}}{f^{2}}\leq\varepsilon_{5}f^{2}|\phi|^{2q+4}+\frac{1}{\varepsilon_{5}}\cdot\frac{|\nabla f|^{2q+4}}{f^{2q+2}}.$
This implies
$\left(A_{\delta}-\sum_{i=1}^{4}\varepsilon_{i}-D\varepsilon_{5}\right)\int_{(M\setminus K)^{+}}f^{2}|\phi|^{2q+4}\leq D\cdot\frac{1}{\varepsilon_{5}}\int_{(M\setminus K)^{+}\cap M_{0}}\frac{|\nabla f|^{2q+4}}{f^{2q+2}}$
$+\left(\frac{(|B|H)^{2}}{\varepsilon_{1}}+\frac{|E_{\delta}+C_{\delta}H^{2}|}{\varepsilon_{2}}+\frac{F^{\frac{2q+4}{3}}}{\varepsilon_{3}}+\frac{|G|^{\frac{q+2}{2}}}{\varepsilon_{4}}\right)\int_{(M\setminus K)^{+}}f^{2}$
The inequality of the statement follows by replacing $f$ with $f^{q+2}$ and letting:
$\beta_{1}$ $=A_{\delta}-\sum_{i=1}^{4}\varepsilon_{i}-D\varepsilon_{5},$
$\beta_{2}$ $=(q+2)^{2q+4}\cdot\frac{D}{\varepsilon_{5}},$
$\beta_{3}$ $=\frac{(|B|H)^{2}}{\varepsilon_{1}}+\frac{|E_{\delta}+C_{\delta}H^{2}|}{\varepsilon_{2}}+\frac{F^{\frac{2q+4}{3}}}{\varepsilon_{3}}+\frac{|G|^{q+2}}{\varepsilon_{4}}.$
Finally, since $\beta_{1}$ is continuous with respect to $\varepsilon_{i},\varepsilon$ and $\tilde{\varepsilon}$, the sign of $\beta_{1}$ at $\varepsilon=\tilde{\varepsilon}=0$ is preserved for sufficiently small values of $\varepsilon_{i},\varepsilon$ and $\tilde{\varepsilon}$. Therefore, it suffices to analyze the sign of $A_{\delta}$ when $\varepsilon=\tilde{\varepsilon}=0$.
We consider the expression:
$A_{\delta}(\varepsilon=\tilde{\varepsilon}=0)=\left(\frac{2}{n}+(2q+1)\right)(1-\delta)-(q+1)^{2}.$
The same computation as in Theorem A.3 shows that
$A_{\delta}>0\quad\text{if and only if}\quad-\delta-\sqrt{\delta^{2}-\delta+\frac{2}{n}(1-\delta)}<q<-\delta+\sqrt{\delta^{2}-\delta+\frac{2}{n}(1-\delta)}.$

### v2

@@ Abstract
We prove that any finite $\delta$-index hypersurface $M$ in ${\mathbb{R}}^{n+1}$ with constant mean curvature must be minimal, provided either of the following conditions holds:
@@
  • the Ricci curvature of $M$ satisfies $\operatorname{Ric}_{M}\geq-\frac{3(1-\delta)}{n-1}|A|^{2}g,$ where $A$ is the second fundamental form and $g$ is the metric on $M.$
@@
We emphasize that no restriction on the dimension is imposed. Moreover, the statement in the second case is new even for finite index hypersurfaces ($\delta=0$).
@@ Introduction
Is a noncompact, complete, stable, constant mean curvature hypersurface of ${\mathbb{R}}^{n+1}$ necessarily minimal?
@@
Do Carmo’s question has been investigated in ${\mathbb{R}}^{n+1}$ and in more general ambient manifolds and in the broader class of finite index constant mean curvature hyperurfaces. A positive answer for $n=3,4$ was obtained independently by [ENR] and Cheng [Che], using the Bonnet-Myers method nowadays also known as the conformal method. More precisely, there is no stable hypersurface with constant mean curvature $H$ ($H$-hypersurface) in a manifold $N$ of dimension $n+1,$ $n=3,4$ provided $H$ is large enough with respect to the sectional curvature of the ambient manifold $N$. In particular, in the Euclidean space this implies nonexistence for any $H\not=0$. Moreover, partial results for a hypersurface $M$ of arbitrary dimension have been obtained by several authors (see for example [AdC] for $M$ with polynomial volume growth , [INS2] for $M$ with subexponential volume growth and [AlCa, DCZh] for $M$ satisfying assumptions on the total curvature, and $n\leq 5$).
@@
A complete, non-compact, finite $\delta$-index, $H$-hypersurface $M$ with
$\operatorname{Ric}_{M}\geq-\frac{3(1-\delta)}{n-1}|A|^{2}g,$
immersed ${\mathbb{R}}^{n+1}$ is a hyperplane.
@@ 1 Definitions and Basic Properties
Throughout this article $M$ will be a complete, noncompact $n$-dimensional hypersurface with constant mean curvature $H$ ($H$-hypersurface) of a Riemannian manifold $N$, $\nu$ its unit normal vector and $A$ its second fundamental form.
@@ 3 Ricci curvature bounded below
$\mathrm{Ric}_{M}\geq-\frac{3(1-\delta)}{n-1}|A|^{2}g.$
@@
$\operatorname{Ric}_{M}\geq-\frac{3(1-\delta)}{n-1}|A|^{2}g.$
@@
In the following subsection we recall some useful definitions and results from criticality theory, pioneered by Murata [Mu] and Pinchover-Tintarev [P1], [P2], [PT]. We also point out that the presentation in [CMMR], together with [PRS], provides a convenient reference for our purposes.
@@ 3.2 The proof of the theorem
$\begin{split}-k\int_{0}^{+\infty}\varphi^{2}\frac{\hat{L}_{V}u}{u}\,ds&+\int_{0}^{+\infty}\varphi^{2}\left(R_{11}+k\hat{V}\right)\,ds+k\int_{0}^{+\infty}\frac{|\nabla u|^{2}}{u^{2}}\\ &\leq(n-1)\int_{0}^{+\infty}(\varphi_{s})^{2}\,ds+k(n-1)\int_{0}^{+\infty}\varphi\varphi_{s}u_{s}u^{-1}\,ds\\ &+\frac{k^{2}(n-1)}{4}\int_{0}^{+\infty}\varphi^{2}u_{s}^{2}u^{-2}\,ds-\int_{0}^{+\infty}\varphi^{2}k(n-2)(\log u)_{ss}\,ds.\end{split}$
@@
$k\int_{0}^{+\infty}\frac{-\hat{L}_{V}u}{u}\varphi^{2}\leq\left(k\left[\frac{(n-1)k}{4}-1\right]+\varepsilon\right)\int_{0}^{+\infty}\varphi^{2}(u_{s})^{2}{u^{-2}}+\left((n-1)+\frac{k^{2}(n-3)^{2}}{4\varepsilon}\right)\int_{0}^{+\infty}(\varphi_{s})^{2}$
@@
Now, take $0\leq\varphi_{R}\leq 1$ such that $\varphi_{R}\equiv 1$ on $[a,R]$, for some fixed $a$, $\varphi_{R}=0$ on $[R,2R]$ and $|(\varphi_{R})_{s}|\leq c/R$, where $c$ is a constant depending on $a.$ Letting $R\to\infty$, one obtains $\hat{L}_{V}u=0$ on $\gamma$. Since the above argument only uses the fact that $u$ is a positive supersolution of $\hat{L}_{V}$, we conclude that every positive supersolution $u$ of $\hat{L}_{V}$ satisfies
@@
Assume by contradiction that $\hat{L}_{V}$ is subcritical. By Theorem 3.5 there exists a positive function $W\in C(M\setminus K)$ such that $\hat{L}_{V}+W\geq 0$. By approximating $W$ from below, we can assume $W\in C^{\infty}(M\setminus K)$. Then, there exists a positive $u\in C^{2,\alpha}_{loc}(M\setminus K)$ such that $\hat{L}_{V}u+Wu\leq 0$, that is $\hat{L}_{V}u\leq-Wu<0$ on the entire $M\setminus K$ and this is impossible.
@@ Appendix A Caccioppoli’s type inequalities
This section is devoted to establishing some Caccioppoli-type inequalities for constant mean curvature $\delta$-stable hypersurfaces. For the reader’s convenience, we present the most significant proofs, which closely follow the arguments in [INS1].
@@
where $R$ is the Riemann curvature tensor of $N$. Furthermore, we assume that the derivative of the curvature tensor is bounded. That is, there exists a constant $K^{\prime}$ such that, for any elements $e_{i},e_{j},e_{k},e_{s},e_{t}$ of a local orthonormal frame, one has
@@
Let $M$ be a noncompact hypersurface immersed in $N$ with constant mean curvature $H$ oriented by a unit normal vector field $\nu$ such that, when $H\not=0,$ then $\vec{H}=H\nu$, with $H>0$. Recall that $\phi$ is the traceless part of the second fundamental form, i.e. $\phi:=A-Hg$, where $g$ is the induced metric on $M$.
@@
[…]
$E_{\delta}$ $=a_{2}(q+1)(q+1+\tilde{\varepsilon})+nK^{2}\left(\frac{2}{n}(1+\varepsilon)+(2q+1)-\tilde{\varepsilon}\right)(1-\delta),$
[…]
for any $\varepsilon,\tilde{\varepsilon}>0$ and
$a_{1}$ $=\frac{n(n-2)}{\sqrt{n(n-1)}},\ a_{2}=\frac{n(K_{2}-K_{1})H}{2}+n(2K_{2}-K_{1}),$
@@
    c=1c=-1, $\varepsilon>0$, $q\in[1,\operatorname{min}\{\alpha_{2},q_{2}\}-\varepsilon]$, and $H^{2}\geq g_{n}(q)$,
where $g_{n},$ $q_{1}$ and $q_{2}$ are defined in the proof.
@@
    The only case with no condition on $|\phi|$ is $A_{\delta}>0$ and $\Delta_{-1}<0$.
    One needs $q\in(\operatorname{max}\{\alpha_{1},q_{1}\},\operatorname{min}\{\alpha_{2},q_{2}\})$ and $H^{2}>g_{n}(q)$, where $g_{n}$ is defined as